OpenStudy (butterflydreamer):

help please :) Two wave fronts off the coast of Sydney meet at the coast. The height at the coast (in cm) of the two wave forms are given by: h1 = 7 + 2 sin t and h2 = 17 + 7cos t . Find the exact value of the initial rate of change of h. ( h = h1 + h2)

OpenStudy (butterflydreamer):

i found that \[h = 24 + \sqrt{53} \sin (t + (\tan^{-1} \frac{ 7 }{ 2 }))\] But i am unsure how to find the initial rate of change.

OpenStudy (fantageplayer):

me neither

OpenStudy (butterflydreamer):

ty.

OpenStudy (phi):

with h= h1 + h2 h1 = 7 + 2 sin t and h2 = 17 + 7cos t I would leave h(t) a sum: h(t) = 24 + 2 sin t + 7 cos t and h' = 2 cos t - 7 sin t

OpenStudy (phi):

by exact value of the initial rate of change, I assume they mean evaluate that expression at t=0 we get h'(0)= 2 (cm/sec ?) not sure of the units

OpenStudy (butterflydreamer):

ohh i see. I totally missed that. Thank you!:)